# Nightjar: Dynamic Adaptive Speculative Decoding for Large Language Models Serving

### v1

@@ Abstract
Speculative decoding (SD) accelerates LLM inference by verifying draft tokens in parallel. However, this method presents a critical trade-off: it improves throughput in low-load, memory-bound systems but degrades performance in high-load, compute-bound environments due to verification overhead. Current SD implementations use a fixed speculative length, failing to adapt to dynamic request rates and creating a significant performance bottleneck in real-world serving scenarios. To overcome this, we propose Nightjar, a novel learning-based algorithm for adaptive speculative inference that adjusts to request load by dynamically selecting the optimal speculative length for different batch sizes and even disabling speculative decoding when it provides no benefit. Experiments show that Nightjar achieves up to 14.8% higher throughput and 20.2% lower latency compared to standard speculative decoding, demonstrating robust efficiency for real-time serving.
Index Terms—  Speculative decoding, Large language models serving, Multi-armed bandit, Dynamic real-time request loads
@@ 1 Introduction
Autoregressive (AR) decoding in large language models (LLMs) is often bottlenecked by memory bandwidth, particularly at small batch sizes, leading to underutilized GPU compute [xia-etal-2024-unlocking, zhang_70_2025]. Speculative decoding (SD) [leviathan2023fast] mitigates this by using a smaller draft model to generate multiple candidate tokens, which the target model then verifies in parallel. This process is lossless in terms of accuracy and increases concurrency. From a roofline perspective (Fig. 1), SD increases arithmetic intensity by verifying multiple tokens per weight, enhancing GPU utilization in the memory-bound regime typical of low batch sizes. Conversely, as the batch size grows and the system becomes compute-bound, the verification overhead of SD negates its benefits, making standard AR decoding more efficient.
@@
Fig. 1: Schematic roofline illustration of autoregressive decoding (w/o SD) and speculative decoding (SD); positions are illustrative.
Fig. 2: Throughput under different request loads and rates. $\gamma$ denotes the speculative length (tokens generated by the draft model per step). Note: Configurations with $\gamma\geq 4$ are omitted at high QPS due to GPU out-of-memory (OOM) errors.
We empirically validate this trade-off [^1] Experiments were conducted on the ShareGPT dataset using vLLM with the DeepSeek-R1-Distill-Qwen-7B model and a 0.5B draft model [alamios_DeepSeek_R1_DRAFT_2024] on an RTX 4090 GPU.. As shown in Fig. 2, SD provides significant speedups at lower request loads. For instance, a speculative length ($\gamma$) of 3 improves throughput by 15.5% at 15 QPS while a speculative length of 2 yielded better improvement than 3 at QPS=25. However, as the load increases, AR decoding surpasses SD’s performance. At high loads (Fig. 2), the overhead becomes detrimental, leading to a performance degradation of up to 30.25% compared to AR. This confirms that SD excels in the memory-bound regime, while its verification overhead becomes a liability when the system is compute-bound.
The core challenge lies in the dynamic and unpredictable nature of real-world LLM serving loads. Consequently, a fixed speculative length is inherently suboptimal, as a configuration tuned for low-load efficiency can degrade service under high traffic. The optimal speculative length is not a static value but a dynamic variable dependent on real-time factors like system load, hardware characteristics, and token acceptance rates [sadhukhan2024magicdec].
While various approaches have sought to optimize speculative decoding, they often fail to address the dynamic and unpredictable nature of real-world LLM serving loads. Initial efforts focused on single-request optimization [wang_opt-tree_2025, zhang_adaeagle_2024, zhang_draft_2024, brown_dynamic_2024, huang_specdec_2024] or static offline batches [su2023synergy, houbanditspec], employing predictive algorithms [zhang_adaeagle_2024] or confidence-based methods [huang_specserve_2025, wang_opt-tree_2025, wu2025tetris], but their static nature makes them fundamentally ill-suited for the fluctuating request rates in live environments. More recent works have explored dynamic speculation length adjustment in serving systems [su2023synergy, liu2024optimizing, huang_specserve_2025, li_adaserve_2025], yet they exhibit significant limitations. For instance, SpecServe [huang_specserve_2025] and TETRIS [wu2025tetris] rely on draft confidence scores, but this can lead to poor performance with large batch sizes and still requires generating at least one token. DSD [liu2024optimizing] optimizes length via linear goodput modeling, but its reliance on past average acceptance rates introduces a critical vulnerability. This dependency on historical data, found in methods like DSD [liu2024optimizing] and SpecServe [huang_specserve_2025], can lead to a “decision deadlock” where disabling speculative decoding permanently halts the collection of new acceptance rate data, locking the system in a suboptimal state. Our exploration-exploitation approach is designed to mitigate this risk. Even the work of BanditSpec [houbanditspec], which first introduced a Multi-Armed Bandit (MAB) paradigm, is constrained by its static design, rendering it incompatible with the continuous batching [kwon2023efficient] of modern serving systems. Furthermore, to the best of our knowledge, existing works largely overlook the “switching cost” incurred when re-enabling speculative decoding from a length of 0, which is the overhead of KV cache reconstruction for the draft model.
Fig. 3: An overview of our adaptive approach with multi-armed bandit (MAB) for dynamic request load.
@@
Therefore, the field urgently needs a truly intelligent, adaptive decision-making mechanism that can perceive load changes in real-time and accurately weigh the trade-offs between benefits and costs. It must be capable of dynamically selecting the optimal speculative strategy for different batch sizes at each decoding step—including decisively disabling it when necessary. To address this core challenge, we propose Nightjar, a contextual bandit approach for speculative length selection in dynamic, real-time LLM services (Fig. 3). A key feature of our approach is its ability to learn the optimal speculative length for different batch sizes at each decoding step, without requiring prior knowledge of request difficulty or model characteristics. Furthermore, we incorporate the switching cost of transitioning from a speculative length of 0 to a positive value, a factor previously overlooked. Implemented on vLLM [kwon2023efficient] and evaluated on three real-world datasets, Nightjar consistently outperforms existing methods, achieving up to 14.8% higher throughput and 20.2% lower latency compared to standard speculative decoding, demonstrating its practical effectiveness for real-time LLM serving.
2 Nightjar Algorithm for Speculative Length Selection
@@ 2.1 Problem Statement
In continual batching [kwon2023efficient] for LLM serving, the batch size can vary at each decoding step. We consider batch sizes $B\in\mathcal{B}=\{1,2,\dots,B_{\max}\}$ and speculative lengths $\gamma\in\{0,1,\dots,\Gamma_{\max}\}$. Following [liu2024optimizing], we evaluate performance using goodput $g_{B,\gamma}$, defined as the number of tokens generated per second, including both accepted proposed tokens and bonus tokens generated during verification.
@@
Our goal is to find a policy that minimizes $R(T)$.
@@ 2.2 Algorithm Design
Contextual MAB can take dynamic request batch size or input sequences as context, such as linear Thompson [agrawal2013thompson], LinUCB [li2010contextual], which have been successfully applied in online prediction settings such as configuration selection [stolcke2023adaptive]. However, they are questionable in dynamic serving environments. We adopt the relatively simple framework ADA-BINGREEDY proposed in [luo2018efficient]. One difference between [luo2018efficient] and our approach is that we develop a request batch mechanism to handle varying request patterns. And we design reward that considers the speculative length switch cost.
Fig. 4: Illustration of the Nightjar algorithm’s three-level hierarchy. Each batch size maintains its own Epoch. Each epoch contains Blocks ($j_{B}$) of exponentially growing duration, which are responsible for the long-term schedule. These blocks are further divided into fixed-size Bins ($b_{B}$), the basic units for controlling the exploration-exploitation trade-off. Each round ($\tau_{B}$) is a decoding step.
As shown in Fig. 4, the Nightjar algorithm adaptively selects the speculative length $\gamma$ for each batch under continuous batching by organizing time into blocks and bins to balance exploration and exploitation.
@@
$\gamma_{t}^{*}=\mathop{\arg\min}\limits_{\gamma\in\{0,1,\dots,\Gamma_{\max}\}}\left\{\frac{1}{\tilde{g}_{B,\gamma}}+\frac{\mathbb{I}(\gamma_{t-1}=0\land\gamma_{t}>0)\cdot c_{\mathrm{prefill}}}{\gamma_{t}}\right\}$ (3)
@@
The first term of Eq. (3), $1/\tilde{g}_{B,\gamma}$, is mathematically dual to maximizing the amortized throughput. The second term functions as a regularization mechanism for the switching cost. The indicator function $\mathbb{I}(\cdot)$ activates only when the system transitions from a disabled state ($\gamma_{t-1}=0$) to an enabled state ($\gamma_{t}>0$), incurring a KV cache reconstruction cost $c_{\mathrm{prefill}}$. By dividing this cost by $\gamma_{t}$, the objective favors configurations that amortize the startup overhead across longer speculative sequences, thereby discouraging short-sighted switching that degrades system latency.
@@
Table 1: Measured switching cost ($c_{\text{prefill}}$) across varying batch sizes and input lengths using a 7B model on an NVIDIA RTX 4090.
@@
We quantify this overhead using the effective skip length, defined as $L_{\text{max}}=\max\{L_{1},\dots,L_{B}\}$, where $L_{i}$ is the number of new tokens for request $i$. Since the draft model verifies the prefix in parallel, the batch latency is determined by the request with the largest lag. The algorithm queries an offline-populated lookup table $c_{\mathrm{prefill}}(L_{\text{max}},B)$ (Table 1) to obtain the precise cost. For example, empirical profiling on a 13B model indicates that $c_{\mathrm{prefill}}$ can range from 5 ms to over 1400 ms depending on $L_{\text{max}}$, making this dynamic cost retrieval essential for robust decision-making.

### v2

@@ Abstract
Speculative decoding (SD) accelerates LLM inference by verifying draft tokens in parallel. However, this method presents a critical trade-off: it improves throughput in low-load, memory-bound systems but degrades performance in high-load, compute-bound environments due to verification overhead. Existing Speculative Decoding strategies typically rely on static speculative lengths, failing to adapt to fluctuating request loads or identify the optimal moment to halt speculation. The cost of restarting speculative inference also remains unquantified. During traffic surges, the marginal utility of speculation diminishes; yet, the draft model’s persistent memory footprint competes for available KV cache. This resource contention limits the maximum batch size, thereby degrading overall system throughput. To overcome this, we propose Nightjar, a novel learning-based algorithm for adaptive speculative inference that adjusts to request load by dynamically selecting the optimal speculative length for different batch sizes. Upon detecting significant request queuing and KV cache shortages, we disable speculative decoding and offload the draft model to the CPU. This reclaims VRAM for the KV cache, thereby facilitating larger batch sizes and maximizing overall system throughput. Experiments show that Nightjar achieves up to 14.8% higher throughput and 20.2% lower latency compared to standard speculative decoding, demonstrating robust efficiency for real-time serving.
@@ 1 Introduction
Large language models (LLMs) are increasingly deployed in real-world applications, where they serve concurrent requests with varying throughput requirements. However, the LLM serving system faces high latency and low throughput due to the sequential nature of autoregressive decoding and the limited memory of the GPU [28].
@@
Speculative decoding (SD) [15, 30] accelerates inference by allowing a draft model to propose multiple candidate tokens, while the target model verifies them in parallel, which is theoretically lossless in terms of accuracy. This reduces the strictly sequential dependency in autoregressive decoding and increases concurrency. From a roofline perspective (Figure 1), under small batch sizes, speculative decoding raises arithmetic intensity by verifying multiple draft tokens per weight, enhancing utilization in the memory-bound regime. As batch size increases, compute demands dominate and the speculative verification overhead pushes the system into the larger compute-bound region, diminishing speculative gains, whereas traditional autoregressive decoding scales more efficiently and increases throughput.
@@
Figure 1: Schematic roofline illustration of autoregressive decoding (w/o SD) and speculative decoding (SD); positions are illustrative.
However, current speculative decoding implementations (e.g., in vLLM [13]) often use a fixed speculative length, which is not optimal for dynamic request loads. The optimal speculative length is a dynamic variable dictated by the interplay of batch size, token acceptance rates, and hardware characteristics. An excessively long speculative length increases the time spent on draft generation and imposes significant verification overhead, whereas a length that is too short fails to fully capitalize on the potential speedup of speculative decoding.
Furthermore, model-based draft models consume both compute resources and precious GPU memory (VRAM), unlike model-free approaches such as N-gram that incur negligible memory footprints. However, model-free methods lack generalizability and are often restricted to specific domains like code generation [20]. In high-traffic scenarios, the VRAM occupied by these draft models directly competes with the KV cache; reclaiming this space for the KV cache would facilitate larger batch sizes and significantly maximize overall system throughput.
Most optimizations for speculative decoding ignore the volatility of real-world serving. Early works targeted single requests [26, 31, 32, 6, 11] or static batches [25, 10], making them unsuitable for fluctuating traffic. Newer dynamic approaches [25, 19, 12, 18] also have flaws. For example, SpecServe [12] and TETRIS [27] rely on draft confidence, which scales poorly with large batch sizes. DSD [19] adjusts length based on historical acceptance rates. This creates a ’deadlock’: if speculation is turned off, data collection stops, trapping the system in a suboptimal state. Even BanditSpec [10], which introduced Multi-Armed Bandits, uses a static design that is incompatible with the continuous batching of modern systems.
While existing works such as Oneiros [17] and eLLM [29] attempt to optimize GPU memory through memory remapping or shared pools, they predominantly focus on general-purpose memory management within or across models. Consequently, they overlook the tight coupling between strategic policy decisions (e.g., speculative length) and memory allocation (e.g., the residency of the draft model) inherent in the speculative decoding paradigm.
@@
Therefore, the field urgently needs a truly intelligent, adaptive decision-making mechanism that can perceive load changes in real-time and accurately weigh the trade-offs between benefits and costs. It must be capable of dynamically selecting the optimal speculative strategy for different batch sizes at each decoding step—including decisively disabling it when necessary. To address this core challenge, we propose Nightjar, a contextual bandit approach for speculative length selection in dynamic, real-time LLM services. A key feature of our approach is its ability to learn the optimal speculative length for different batch sizes at each decoding step, without requiring prior knowledge of request difficulty or model characteristics. Furthermore, we incorporate the switching cost of transitioning from a speculative length of 0 to a positive value, a factor previously overlooked. By dynamically offloading the draft model upon disabling speculative inference, we reclaim substantial VRAM for the KV cache, thereby facilitating larger batch sizes and maximizing overall system throughput. Simultaneously, when a reduction in request load is detected, we proactively and asynchronously reload the draft model to re-enable speculative decoding, ensuring sustained performance gains. Implemented on vLLM [13] and evaluated on three real-world datasets, Nightjar consistently outperforms existing methods, achieving up to 14.8% higher throughput and 20.2% lower latency compared to standard speculative decoding, demonstrating its practical effectiveness for real-time LLM serving.
2 Motivation
Although modern LLM serving systems, e.g., vLLM [13], have been widely adopted in production, challenges remain in achieving high throughput and low latency with speculative decoding.
2.1 Dynamic Speculative Decoding Length Selection
Figure 2: Throughput under different request loads and rates. $\gamma$ denotes the speculative length (tokens generated by the draft model per step). Note: Configurations with $\gamma\geq 4$ are omitted at high QPS due to GPU out-of-memory (OOM) errors.
In speculative decoding, the draft model generates $\gamma$ tokens, which are subsequently verified by the target model in parallel. Consequently, the total latency per iteration is the sum of the draft model’s generation time and the target model’s verification time. The actual number of tokens produced per step equals the number of accepted tokens ($n$) plus one bonus token. While increasing the speculative length $\gamma$ can potentially raise the number of accepted tokens, it also incurs a higher temporal cost. If the acceptance rate (the ratio of accepted tokens to $\gamma$) is low, the added overhead may outweigh the gains, ultimately leading to a decrease in overall throughput.
We empirically validate this trade-off [^1] Experiments were conducted on the ShareGPT dataset using vLLM with the DeepSeek-R1-Distill-Qwen-7B model and a 0.5B draft model [2] on an RTX 4090 GPU.. As shown in Figure 2, SD provides significant speedups at lower request loads. For instance, a speculative length ($\gamma$) of 3 improves throughput by 15.5% at 15 QPS while a speculative length of 2 yields better improvement than 3 at QPS=25. However, as the load increases, Autoregressive (AR) decoding surpasses SD’s performance. At high loads (Figure 2), the overhead becomes detrimental, leading to a performance degradation of up to 30.25% compared to AR. This confirms that SD excels in the memory-bound regime, while its verification overhead becomes a liability when the system is compute-bound.
The core challenge lies in the dynamic and unpredictable nature of real-world LLM serving loads. Consequently, a fixed speculative length is inherently suboptimal, as a configuration tuned for low-load efficiency can degrade service under high traffic. The optimal speculative length is not a static value but a dynamic variable dependent on real-time factors like system load, hardware characteristics, and token acceptance rates [23].
Figure 3: Memory occupation for different draft methods.
2.2 Memory underutilization
TABLE I: KV Cache memory occupation for different models.
Context Length Vicuna-13B (MHA) DeepSeek-R1-Distill-Qwen-7B
1,024 (1K) 800 MiB 56 MiB
4,096 (4K) 3,200 MiB (3.12 GiB) 224 MiB
8,192 (8K) 6,400 MiB (6.25 GiB) 448 MiB
32,768 (32K) 25,600 MiB (25.00 GiB) 1,792 MiB (1.75 GiB)
As evidenced by Table I, the memory footprint of the KV cache is not only substantial but scales aggressively with context length, transitioning from a negligible overhead to a primary consumer of GPU resources. This expansion creates a zero-sum competition within the finite VRAM budget: as the KV cache encroaches upon the available memory, it directly constrains the headroom for resident draft model weights. Such resource contention highlights a critical bottleneck in static memory partitioning, where the burgeoning cache demands can eventually marginalize auxiliary models and degrade overall system throughput.
As shown in Figure 3, draft models (Ngram, DeepSeek-R1-DRAFT-Qwen2.5-0.5B [2]) occupy varying GPU memory, impacting KV cache capacity. While their weights are smaller than the main model, they constrain memory under high loads. For instance, unloading a 1.4GB draft model (e.g., 14B Multi-Token Prediction (MTP) for DeepSeek 671B [9]) frees space for  49k additional tokens in KV cache (calculated via MLA [9]). Current systems (e.g., vLLM [13]) retain draft models even when SD is disabled, wasting memory.
We identify that memory inefficiency arises from the isolation of weight and KV cache spaces in existing systems. This isolation hinders the dynamic redistribution of memory resources based on workload demands, leading to underutilized activation memory, which exacerbates KV cache space pressure and ultimately constrains overall system performance.
3 Nightjar Design
In this section, we first introduce the problem statement and the key notations used in Nightjar. Then, we present the algorithm design and the elastic memory management mechanism for Speculative Decoding.
3.1 Overview
Figure 4: An overview of our adaptive approach with multi-armed bandit (MAB) for dynamic request load.
Nightjar consists of two core components: 1) Speculative Length Selection: A contextual multi-armed bandit (MAB) algorithm that selects the optimal speculative length for each batch size in continual batching. If the speculative length is selected to be 0, the draft model will not be used and the system will switch to the standard autoregressive decoding. Otherwise, the draft model will be used to generate the speculative length tokens, and the target model will verify them in parallel. 2) Elastic Memory Management: A mechanism that dynamically allocates memory for the draft model. If the draft model is not used, and request rate is high, and there is not enough GPU memory for KV cache to allocate, we free the draft model in the GPU and make room for the KV cache. When the request rate is low, we load the draft model back into the GPU to accelerate the generation, as illustrated in Figure 4.
4 Multi-Armed Bandit for Speculative Length Selection
In this section, we introduce the multi-armed bandit (MAB) for speculative length selection.
@@ 4.1 Problem Statement
In continual batching [13] for LLM serving, the batch size can vary at each decoding step. We consider batch sizes $B\in\mathcal{B}=\{1,2,\dots,B_{\max}\}$ and speculative lengths $\gamma\in\{0,1,\dots,\Gamma_{\max}\}$. Following [19], we evaluate performance using goodput $g_{B,\gamma}$, defined as the number of tokens generated per second, including both accepted proposed tokens and bonus tokens generated during verification.
@@
Our goal is to find a policy that minimizes $R(T)$. We summarize the key notations and their descriptions in Table II.
TABLE II: Summary of Key Notations used in Nightjar.
Symbol Description
System Parameters & Cost Model
$B,B_{\max}$ Current and maximum batch size, where $B\in\{1,\dots,B_{\max}\}$.
$\gamma,\Gamma_{\max}$ Current and maximum speculative length (draft tokens).
$c_{\mathrm{prefill}}$ Latency overhead for reconstructing the KV cache (switching cost).
$L_{\max}$ Effective skip length, defined as $\max_{i\in B}L_{i}$ (max lag in a batch).
$C_{\text{switch}}$ Upper bound constant for the switching cost.
Elastic Memory Management
$M_{\text{orig}},M_{\text{scale}}$ Baseline KV cache capacity and expanded capacity after offloading.
$S_{\text{draft}}$ Memory footprint required to store the draft model weights.
$N_{\text{free}}$ Number of currently available physical KV cache blocks.
$\tau_{\text{low}}$ Critical scarcity threshold triggering memory expansion.
$T_{\text{persist}}$ Temporal window size for validating low-memory states.
$|\mathcal{Q}_{\text{wait}}|$ Cardinality of the request waiting queue.
MAB Hierarchical Structure
$t$ Global decoding time step index.
$j_{B}$ Block index for batch size $B$, governing long-term scheduling scales.
$b_{B}$ Bin index within block $j_{B}$, controlling exploration probability ($p\propto 1/\sqrt{b_{B}}$).
$\tau_{B}$ Round counter within bin $b_{B}$.
$H_{B}$ Duration of the current block, growing exponentially as $2^{j_{B}-1}$.
Performance Metrics & Regret Analysis
$g_{B,\gamma},\tilde{g}_{B,\gamma}$ Theoretical goodput and its empirical mean estimate.
$\gamma^{*}_{t}$ Optimal speculative length at time $t$ that minimizes the objective.
$r_{t}(\gamma)$ Realized reward (observed goodput) at step $t$ for arm $\gamma$.
$N_{t}(\gamma)$ Number of times arm $\gamma$ has been selected up to time $t$.
$\beta_{t}(\gamma)$ Confidence radius for the goodput estimate at time $t$.
$R(T)$ Cumulative regret over horizon $T$ relative to the optimal policy.
$\delta$ Confidence parameter, where estimates hold with prob. $1-\delta$.
@@ 4.2 Algorithm Design
Contextual MAB can take dynamic request batch size or input sequences as context, such as linear Thompson [1], LinUCB [16], which have been successfully applied in online prediction settings such as configuration selection [24]. However, they are questionable in dynamic serving environments. We adopt the relatively simple framework ADA-BINGREEDY proposed in [22]. One difference between [22] and our approach is that we develop a request batch mechanism to handle varying request patterns. And we design reward that considers the speculative length switch cost.
Figure 5: Illustration of the Nightjar algorithm’s three-level hierarchy. Each batch size maintains its own Epoch. Each epoch contains Blocks ($j_{B}$) of exponentially growing duration, which are responsible for the long-term schedule. These blocks are further divided into fixed-size Bins ($b_{B}$), the basic units for controlling the exploration-exploitation trade-off. Each round ($\tau_{B}$) is a decoding step.
As shown in Figure 5, the Nightjar algorithm adaptively selects the speculative length $\gamma$ for each batch under continuous batching by organizing time into blocks and bins to balance exploration and exploitation.
@@
$\gamma_{t}^{*}=\mathop{\arg\min}\limits_{\gamma\in\{0,1,\dots,\Gamma_{\max}\}}\left\{\frac{1}{\tilde{g}_{B,\gamma}}+\frac{\mathbb{I}(\gamma_{t-1}=0\land\gamma>0)\cdot c_{\mathrm{prefill}}}{\gamma}\right\}$ (3)
@@
The first term of Eq. (3), $1/\tilde{g}_{B,\gamma}$, is mathematically dual to maximizing the amortized throughput. The second term functions as a regularization mechanism for the switching cost. The indicator function $\mathbb{I}(\cdot)$ activates only when the system transitions from a disabled state ($\gamma_{t-1}=0$) to an enabled state ($\gamma>0$), incurring a KV cache reconstruction cost $c_{\mathrm{prefill}}$. By dividing this cost by $\gamma$, the objective favors configurations that amortize the startup overhead across longer speculative sequences, thereby discouraging short-sighted switching that degrades system latency.
@@
TABLE III: Measured switching cost ($c_{\text{prefill}}$) across varying batch sizes and input lengths using a 7B model on an NVIDIA RTX 4090.
@@
We quantify this overhead using the effective skip length, defined as $L_{\text{max}}=\max\{L_{1},\dots,L_{B}\}$, where $L_{i}$ is the number of new tokens for request $i$. Since the draft model computes the tokens in parallel, the batch latency is determined by the request with the largest lag. The algorithm queries an offline-populated lookup table $c_{\mathrm{prefill}}(L_{\text{max}},B)$ (Table III) to obtain the precise cost. For example, empirical profiling on a DeepSeek-R1-Distill-Qwen-7B model target model and 0.5B draft model on RTX 4090 indicates that $c_{\mathrm{prefill}}$ can range from 17.87 ms to over 102.03 ms depending on $L_{\text{max}}$, making this dynamic cost retrieval essential for robust decision-making. To construct this lookup table, we performed offline profiling. We measured latencies across a grid of input lengths and geometrically increasing batch sizes (e.g., 2, 4, 8…). Specifically, we recorded the prefill time for the standalone target model ($T_{\text{base}}$) and the prefill time with speculative decoding enabled ($T_{\text{SD}}$). The switching cost $c_{\text{prefill}}$ is derived from the difference ($T_{\text{SD}}-T_{\text{base}}$).